Find the cube root of each of the following numbers by prime factorisation method.

(i) 64

 

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student
(i) 64

Solution:

64 = 2×2×2×2×2×2

By grouping the factors in triplets of equal factors, 64 = (2×2×2)×(2×2×2)

Here, 64 can be grouped into triplets of equal factors,

∴ 64 = 2×2 = 4

Hence, 4 is cube root of 64.




Thanks